# RadarPLM: Adapting Pre-trained Language Models for Marine Radar Target Detection by Selective Fine-tuning

### v4

@@ IV-D1 Backbone Architectures
IV-D2 Fine-tuning Strategies
@@ IV-D3 Loss Function
IV-E Analysis of the Effect of Hyperparameters on Experimental Results
@@ V Conclusion
In this study, we propose RadarPLM, a novel fine-tuning framework to adapt PLMs for marine radar target detection. We design a lightweight fine-tuning module for lightweight adaptation and a preference-aware loss for selective optimization. Extensive experimental results demonstrate that RadarPLM significantly outperforms state-of-the-art baselines on low SCR datasets and scarce training data settings. These results highlight the remarkable potential of PLMs as a general optimization solver for radar signal processing. Looking ahead, future research will focus on leveraging larger open-source PLMs, and incorporating model compression techniques such as quantization, pruning, and knowledge distillation to further enhance RadarPLM’s scalability and deployment efficiency.

### v5

@@ IV-E1 Backbone Architectures
IV-E2 Fine-tuning Strategies
@@ IV-E3 Loss Function
IV-F Analysis of the Effect of Hyperparameters on Experimental Results
@@ V Conclusion
We propose a fine-tuning framework RadarPLM that adapts the GPT2, an advanced open-source pre-trained language model, for marine radar target detection. To ensure lightweight and robust adaptation, we devise a novel lightweight adaptation module, preserving the general knowledge of the GPT2 model while significantly reducing training time and fine-tuning overhead. To mitigate overfitting on low-SCR radar signal, we develop a selective training strategy to selectively optimize different feature patches based on their online-evaluated learning values, effectively improving model robustness. Extensive experiments on real-world marine radar datasets demonstrate that RadarPLM consistently outperforms SOTA baselines, especially under low SCR conditions, achieving an average detection rate improvement exceeding 6% over the strongest competitors. Notably, under small-sample training conditions, RadarPLM surpasses the existing methods by nearly 20% at most, owing to the effective transfer of the universal knowledge in GPT-2 to the marine radar signal processing task.